MILLIMETER WAVE SPECTROSCOPY AND MQDT CALCULATIONS OF HIGH RYDBERG STATES OF KRYPTON

\underline{MARTIN SCHÄFER} and FRÉDÉRIC MERKT, ETH Zürich, Laboratorium für Physikalische Chemie, ETH-Hönggerberg HCI, CH-8093 Zürich, Switzerland.

A phase-stabilised backward wave oscillator (BWO) in the 260–380 GHz range was combined with a VUV laser system to record high-resolution spectra of high-n Rydberg states of krypton. Krypton atoms were excited into np (n=58,60) Rydberg states via the 4d[1/2] (J = 1) state using VUV and visible laser photons. Millimeter wave transitions between np and ns or nd Rydberg states were detected by pulsed field-ionization, at sub-MHz resolution and with mass selection. Using this excitation scheme, very accurate relative energies of fine and hyperfine structure levels of (n + 2)s and nd (n=68–74) Rydberg states of ^{84}Kr and ^{83}Kr were obtained. In this region, s-d interactions are observable for ^{83}Kr due to the hyperfine interaction. A multichannel quantum defect theory (MQDT) treatment of the hyperfine structure^{a} was used to analyze the millimeter wave data in combination with recent high-resolution VUV laser data^{a} and the available data from the literature; improved MQDT parameters and hyperfine structure data of the ^{2}P ground electronic state of ^{83}Kr^{+} were obtained.

aH. J. Wörner, U. Hollenstein, and F. Merkt, Phys. Rev. A, 68, 032510 (2003).